let ABC be an equilateral triangle inscribed in a circle x2+y2=a2. suppose perpendiculars from A, B, c to the major axis of the ellipse (x2/a2)+(y2/b2)=1, (a>b) mets ellipse respectively at P, Q, R so that P, Q, R lie on the same sideof the major axis as A,B,C resp. Then prove that Normals to ellipse at P, Q, R are concurrent.

Grade:12

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

I will just give u a hint:

The equation of normal to the ellipse (x2/a2) + (y2/b2) = 1
At the point (a cosθ,bsinθ) is
ax Sec θ - by Cosec θ = a2 - b2

 

Concurrent lines or line segments occur when 3 or more lines meet at the same point.

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi